OpenStudy (ccastudentgrade10):

PLEASE HELP ME |-4b - 8| + |-1 - b^2| + 2b^3; b = -2 please explain steps?

OpenStudy (bunnykitten02):

i ca t really help you but if you go to brainly.com they can help, thats were my sister got her answer and her quesion was pretty close to this

OpenStudy (ccastudentgrade10):

okay ill try that

OpenStudy (anonymous):

substitute -2 for every b. Take it one piece at a time. Do you know how to do this part |-4(-2) - 8| =

OpenStudy (ccastudentgrade10):

why do you substitute -2 for every b? (i have to explain my steps) @peachpi

OpenStudy (ccastudentgrade10):

oh wait duh it makes so much sense now! but how do i work in the absolute value? do i just ignore it?

OpenStudy (anonymous):

sub the 2 because that's what they're asking you to do. don't ignore the absolute value signs. Make what you end up with positive, then drop them.

OpenStudy (anonymous):

|-4(-2) - 8| = |8 - 8| = |0| = 0

OpenStudy (ccastudentgrade10):

ohhhh okay thank you!!

OpenStudy (anonymous):

try this part |-1 - b^2|

OpenStudy (ccastudentgrade10):

so it would be |-1 - (-2)^2 |-1 - -4| |-1 + 4| |3| =3

OpenStudy (ccastudentgrade10):

@peachpi

OpenStudy (ccastudentgrade10):

for that part

OpenStudy (anonymous):

close. When you do (-2)² you get should have gotten 4 because a negative times a negative is a positive

OpenStudy (ccastudentgrade10):

oh right! okay thank you! when i get my answer ill type it here

OpenStudy (anonymous):

ok

OpenStudy (ccastudentgrade10):

is it -11? @peachpi

OpenStudy (anonymous):

yes -11

OpenStudy (ccastudentgrade10):

thank you!

OpenStudy (anonymous):

you're welcome